OpenStudy (anonymous):

NEED HELP!!! WILL GIVE MEDAL!!! What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3

OpenStudy (anonymous):

answer choices: f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ash2326):

@cambam1323 did you try plugging in x and finding f(x)?

OpenStudy (anonymous):

no i have no idea how to find these problems. will that work by plugging in for x and i get fx on the right equation? its just a matter of plugging x?

OpenStudy (ash2326):

We are given this x f(x) 1 -4 3 -1 4 3 so our answer is the one that will hold true for all of these relations. You can start by putting x=1 and check f(x) for all the options, the ones which give f(x)=-4 may be the answer but the ones which are anything but -4, we can certainly ignore them. Do you follow? can you try?

OpenStudy (anonymous):

yes hang on let me give it a try. i know already that the first two wont work because i tried them.

OpenStudy (ash2326):

yes, take your time.

OpenStudy (anonymous):

oops each x-1 is supposed to be the ^x-1 of (3)... the first two still don't work though.

OpenStudy (ash2326):

could you check second option again?

OpenStudy (anonymous):

sure! i was thinking the last one but it doesnt hurt to double check! :)

OpenStudy (ash2326):

yes, good ;)

OpenStudy (anonymous):

so it is the last one? i just double checked 2 and it didnt work.

OpenStudy (anonymous):

but the last one does.

OpenStudy (ash2326):

I believe second is also a possible option based on x=1 and f(x)=-4 \[f(x)=4 (x-1)^3-4\] Let's plugin x=1 \[f(x)=4 (1-1)^3-4\] \[f(x)=4\times 0 -4\] \[f(x)=0-4\longrightarrow -4\]

OpenStudy (anonymous):

\[f(x)=4(3)^{x-1}-4 \] is how option two looks.

OpenStudy (ash2326):

oh I am sorry, then you are correct

OpenStudy (anonymous):

okay yay! thank you for helping me!!! :)

OpenStudy (ash2326):

no problem. Apologies for the confusion :/

OpenStudy (anonymous):

its okay that happens alot in math and technology for that matter haha

OpenStudy (ash2326):

ha ha :D